OpenStudy (anonymous):

pls can someone help me with this; Determine the following integrals by using trigonometric substitution explicitly indefinite integral 10/x^2+9dx

thomaster (thomaster):

\(\Large\int{\frac{10}{x^2}+9dx}\) or \(\Large\int{\frac{10}{x^2+9}dx}~~?\)

OpenStudy (anonymous):

@thomaster, since we're using a trig sub here, it's likely the second.

thomaster (thomaster):

@SithsAndGiggles hehe english math terms are not my best point :P

OpenStudy (anonymous):

ya, the second one

OpenStudy (anonymous):

Let \(x=3\tan u\). Then replace \(dx\) with an expression containing \(du\), and the rest should follow through nicely.

OpenStudy (loser66):

got it? friend

OpenStudy (anonymous):

na, i'm kinda stuck at the part where i supposed to substitute. should i substitute 3tanu where x^2 is?

OpenStudy (loser66):

nope, it is let x = 3tan u x^2 = 9tan^2 u +9 both sides x^2 +9 = 9tan^2 u + 9 = 9(tan^2 u +1) = 9 sec^2 u

OpenStudy (anonymous):

oh ok, just got it!

OpenStudy (loser66):

good

OpenStudy (anonymous):

my answer is 10/3u+c.

OpenStudy (loser66):

don't forget change the limits

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

ah, ok. thnx a lot

OpenStudy (loser66):

I don't know, let me try.

OpenStudy (loser66):

oh, you don't have limits, sorry,

OpenStudy (loser66):

but it not that result, you have trig on the result

OpenStudy (loser66):

\[\frac{10}{9sec^2 u}= \frac {10cos^2u}{9}\]

OpenStudy (loser66):

and it's quite easy to take integral

OpenStudy (loser66):

got what I mean?

OpenStudy (anonymous):

ya :)

OpenStudy (loser66):

ok, good

OpenStudy (anonymous):

but i think u mean sec^2u instead of cos ^2u

OpenStudy (loser66):

nope, \[sec u = \frac{1}{cos u}---> \frac{1}{sec u} = cos u\]

OpenStudy (loser66):

right?